OpenStudy (anonymous):

Write the equation y - 4 = -2(x - 5) in slope-intercept form. A. y = -2x + 14 B. y = -2x + 5 C. y = -2x + 9 D. y = -2x + 6

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Ok, simply expand so and move stuff about to get it into the form \(y = mx +c\) where m is the slope and c is the intercept. First step could be to expand the right hand side. Can you do that?

OpenStudy (anonymous):

You got me confused.

OpenStudy (anonymous):

so the right hand side reads \[-2(x-5)\] right? How would you expand that so that there is no brackets?

OpenStudy (anonymous):

-2x--10?

OpenStudy (dwelsher):

put it in the slope intercept equation . this an important equation that your need to always remember-y=mx+b. where m = the slope and b equals the y intercept

OpenStudy (dwelsher):

y= mx+b that is

OpenStudy (anonymous):

Exactly! But we can simplify further to get rid of the double negative \[-2x+10\]Does that make sense? Now we have \[y-4=-2x+10\]

OpenStudy (anonymous):

Ok. Then do we divide the -2x+10 by 4? I don't really understand all of this at all..

OpenStudy (anonymous):

Remember we are trying to get the equation in the form \[y = mx +c\]so we could add something to both sides to get rid of the 4 on the Left hand side. \[y-4 = -2x+10\] Now add 4 to both sides\[y-4 +4 = - 2x + 10 + 4\]What does this end up as?

OpenStudy (anonymous):

y= -2x+14?

OpenStudy (anonymous):

Great job!

OpenStudy (anonymous):

Thanks

OpenStudy (dwelsher):

remember that the slope equals y/x=m